Five cards are drawn from a standard 52 card declk. What is the probibility that the first 2 cards are red?
I put the formula as 26! / (26!-2!)*2!
But I'm not sure its right.
Please help, thanks.

Answer by stanbon(75887)

Five cards are drawn from a standard 52 card declk. What is the probibility that the first 2 cards are red?
I put the formula as 26! / (26!-2!)*2!
---------------------------------------------
# of ways to get 2 red: 26C2
-----
# of ways to get 3 black: 26C3
--------
# of 5 card hands: 52C5
----------------------------------
P(2 red then 3 black) = 26C2*26C3/52C5 = 845,000/2598960 = 0.3251
====================
Cheers,
Stan H.
===================